Chi-Square Testing 

A survey was carried out in a firm of the smoking habits of men and women employees with the following results:

       

Men Women

Smokers 48 27

Non-smokers 58 57

 

It is required to test whether, at the 5% level of significance, the survey reveals any difference in the smoking habits of men and women.


1
Expert's answer
2021-08-18T14:07:01-0400

Expected values:

"\begin{matrix}
 & Men&Women \\
 Smokers & 41.84 & 33.16\\
 Non-smokers & 64.16 & 50.84
\end{matrix}"

Test statistic: "\chi^2=\frac{(48-41.84)^2}{41.84}+\frac{(27-33.16)^2}{33.16}+\frac{(58-64.16)^2}{64.16}+\frac{(57-50.84)^2}{50.84}=3.387."

Degrees of freedom: "df=(2-1)(2-1)=1."

P-value: "p=P(\chi^2>3.387)=0.0657."

Since the p-value is greater than 0.05, fail to reject the null hypothesis.

The survey did not reveal any difference in the smoking habits of men and women.